Electrodynamics A (PHY 5346) Fall 2016 Classwork and Homework

Every exercise counts 10 points unless stated differently.

Set 1:

1. Exercise E.2: A non-relativistic distance measurement. Homework, due 9/13/2016 before class.

2. Exercise E.4: Non-relativistic Galilei transformation. Homework, due 9/13/2016 before class.

3. Exercise E.5: x_{α}x^{α} contractions. Homework, due 9/13/2016 before class.

4. Time and relativistic distance measurements. Classwork, due 9/1/2016 in class. A Cesium clock counts 2,757,789,531,312 cycles (starting at 0). Find the elapsed time in seconds (you may round to the nearest integer number)? An observer O_{1} is located in an inertial system and flashes at times 1, 2, 3 … [s] light signals towards another observer O_{2}, who reflects them back with a mirror. Assume that the returned signals are received by O_{1} at the following times:

(1) 1.002, 2.002, 3.002 … [s],

(2) 1.002, 2.004, 3.006 … [s],

(3) 1.002, 2.004, 3.008 … [s]. Determine for each case whether the data are consistent with assuming that O_{2} is also in an inertial system. If this is the case, write down the equation for the distance of O_{2} as function of the time as seen by O_{1}. Approximate the speed of light by 300, 000 [km/s], but perform all calculations with a precision of at least four digits.

Set 2:

5. Exercise E.6: Time in Minkowski space. Homework, due 9/15/2016 before class.

6. Exercise E.8: Euclidean and hyperbolic rotations. Homework, due 9/15/2016 before class.